• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Jack (for Heidi)

    Log entry time 18:27:20 on December 18, 2007

    Entry number 221858

    keyword=Left Q1 He level readback

    The Left Q1 He level readback had an erratic offset in the cryo crate. If it is wired similar to the Right Q1 (see halog entry 213020) it will take a few hours to correct. It was set to what seems a more stable configuration for the moment. It will be investigated further thursday. If Left Q1 trips on a spurious low He level just reset the power supply. Sorry for the inconvenience.